OpenStudy (anonymous):

ok so please tell me this is right m-13=-8 ---- 2 m=-6

jimthompson5910 (jim_thompson5910):

is this \[\Large \frac{m-13}{2}=-8\] ???

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\Large \frac{m-13}{2}=-8\] \[\Large m-13=-8*2\] \[\Large m-13=-16\] \[\Large m=-16+13\] \[\Large m=-3\]

OpenStudy (anonymous):

ooooooooo i did 10 sorry i was looking at another problem

jimthompson5910 (jim_thompson5910):

ah ok

OpenStudy (anonymous):

ok so k+10=5 ____ -2 k=20

jimthompson5910 (jim_thompson5910):

not quite, multiply both sides by -2, what do you get?

OpenStudy (anonymous):

-10

jimthompson5910 (jim_thompson5910):

on the right side, what about the left?

OpenStudy (anonymous):

-20

jimthompson5910 (jim_thompson5910):

no, multiplying the left side by -2 gives you k+10

jimthompson5910 (jim_thompson5910):

this is because the -2 terms will cancel

jimthompson5910 (jim_thompson5910):

on the left side

OpenStudy (anonymous):

could you draw that out please

jimthompson5910 (jim_thompson5910):

\[\Large \frac{k+10}{-2}=5\] \[\Large k+10=5(-2)\] \[\Large k+10=-10\] \[\Large k=-10-10\] \[\Large k=-20\]

OpenStudy (anonymous):

didnt i have -20

jimthompson5910 (jim_thompson5910):

you had 20

jimthompson5910 (jim_thompson5910):

so you were close